ganeshie8 (ganeshie8):

Evaluate \[\large \int \limits_0^{\infty}\frac{e^{-x}}{\sqrt{x}} \, dx = \sqrt{\pi} \] @Marki

OpenStudy (turingtest):

i found the answer online but i don't get one step, so I won't bring it up :P

OpenStudy (anonymous):

ok so far i got 2 methods :)

OpenStudy (anonymous):

\[\int\limits_{ }^{ } \frac{e^{-x}}{\sqrt{x}}dx\]\[2\sqrt{x}e^{-x}+2\int\limits_{ }^{ }\sqrt{x}e^{-x}~dx\](differentiated e^-x, integ. by parts) \[2\sqrt{x}e^{-x}+2\int\limits_{ }^{ }\sqrt{x}e^{-x}~dx\]u sub, solving the second integral. u=sqrt{x} -x=-u^2 du= 1/(2sqrt{x}) dx >> 2sqrt{x}du=dx >> 2u du=dx 2 * integ. 2u^2 e^(-2u) du integration by parts twice to reduce the u^2.

hartnn (hartnn):

Laplace :3

OpenStudy (anonymous):

my method, if I haven't made a mistake, is not hard...

ganeshie8 (ganeshie8):

Oh nice i have searched a bit in MSE but couldnt find this :O please provide the link @TuringTest :)

ganeshie8 (ganeshie8):

Interesting xD f(t) = 1/sqrt(t) and we want L(1) @hartnn ? @((*-*)) im still going thru your method

OpenStudy (anonymous):

take your time

hartnn (hartnn):

yes ganesh, thats the easiest way :P

OpenStudy (turingtest):

http://math.stackexchange.com/questions/215352/why-is-gamma-left-frac12-right-sqrt-pi (I am looking at the one using spherical coordinates) I don't get why the bounds change on the u-sub, and why the u-sub is correct shouldn't it lead to 4x in the exponent?

OpenStudy (michele_laino):

I think, it is necessary integrate by parts, first, as below: \[e ^{-x}*2\sqrt{x}|_{0}^{+\infty}+2\int\limits_{0}^{+\infty}e ^{-x}\sqrt{x}dx=2\Gamma(3/2)=2*\frac{ \sqrt{\pi} }{ 2 }=\sqrt{\pi}\]

OpenStudy (anonymous):

lets take indefinite integral as refrence \(\large \int \limits_{-\infty}^{\infty} {e^{-x^2}} \, dx \\=\large 2\int \limits_0^{\infty} {e^{-x^2}} \, dx \) x= sqrt t so we would have \(\large 2\int \limits_0^{\infty} {e^{-x^2}} \, dx =\large 2\int \limits_0^{\infty} \frac{1}{2} {e^{-t}} t^{-\frac{1}{2} } \, dx = \Gamma(1/2) \) now lets proof \(\large \int \limits_{-\infty}^{\infty} {e^{-x^2}} \, dx \) there is several way chose from here to make me explain :P http://en.wikipedia.org/wiki/Gaussian_integral

OpenStudy (anonymous):

u know that definite integral from -infinity to infinity give -sqrt pi now only chose any proof in wiki link so i would explain ,till then i'll post next method

OpenStudy (anonymous):

sorry indefinite*

ganeshie8 (ganeshie8):

that looks very neat xD I can work \(\int_0^{\infty} e^{-x^2}dx\) using polar coordinates.. lm going through that mse link provided by TT surprisingly nobody used the obvious laplace transform there.. is it easy to find the laplace transform of 1/sqrt(t) ?

OpenStudy (turingtest):

@hartnn mentioned that laplace was the easy way

ganeshie8 (ganeshie8):

@Michele_Laino how did you get \(\Gamma (3/2) = \sqrt{\pi}/2\) only if we can get some integer after doing by parts n times xD

OpenStudy (anonymous):

hmm i never used or liked laplace :P my second method is using gama function properties im just writing my notes now , will provide u with it quick

OpenStudy (anonymous):

I was like, hey guys, how about an elementary calculus?

OpenStudy (turingtest):

@ganeshie8 what I don't understand is why it changes to\(\int_{-\infty}^\infty e^{-x^2}dx\) and *not* \(\int_{0}^\infty e^{-x^2}dx\) could somebody explain that to me? I am dumb

OpenStudy (michele_laino):

since for a property of the Gamma function, we can write: \[\Gamma (z+1)=z \Gamma (z)\] so: \[\Gamma(3/2)=\Gamma (1+1/2)=\frac{ 1 }{ 2 }\Gamma \left( \frac{ 1 }{ 2 } \right)\] @ganeshie8

OpenStudy (anonymous):

well know that \(\Gamma (p)=p\Gamma (p-1)\) so 3/2=1+1/2 \(\Gamma (3/2)=3/2\Gamma (1/2)\) hmm how did u got sqrt pi /2 ?

OpenStudy (anonymous):

oh wait my bad :P its p-1 gamma (p-1)

OpenStudy (anonymous):

then its true :O

OpenStudy (turingtest):

my questions got ignored :P fine, I'll figure it out myself

ganeshie8 (ganeshie8):

hey no im still thinking.. it has to do with e^(-x^2) being an converging even function i think

OpenStudy (turingtest):

hm i didnt think of that

ganeshie8 (ganeshie8):

\[2\int\limits_0^{\infty} e^{-x^2} dx =\int\limits_{-\infty }^{\infty} e^{-x^2} dx \]

OpenStudy (turingtest):

specifically\[\int_0^\infty e^{-x}d(2\sqrt x)=\int_{-\infty}^\infty e^{-u^2}du\]

OpenStudy (turingtest):

ok first of all where did the 2 go?

OpenStudy (michele_laino):

sorry @Marki I already deleted my post

OpenStudy (anonymous):

well ,this integral is the base of all integrals in complex :P i wuld say its like number 1 in multiplying operation and @Michele_Laino i note that thank u :) , how ever @TuringTest